# Loops, Holonomy and Signature

### v3

@@ 5.1. Main lemma
If $\gamma\in\Omega$ is not trivial in ${\mathcal{E}}^{\,psi,ret}(\mathbb{R}^{n})$, then there are compactly supported one forms $\omega_{1},\ldots,\omega_{q}$ such that
@@
Since $SL(2,\,\mathbb{R})$ is simple, then Theorem 2.1 applies, that is, there is a connection $A$ in the bundle $\mathbb{R}^{n}\times SL(2,\mathbb{R})$, such that its holonomy $H_{A}(\gamma)$ is not trivial. This can also be deduced from Tlas result [Tl] and Proposition 1.1 as follows: since every whisker in the piecewise smooth class is a thin loop (see Remark $4$ in [Me]), by Proposition 1.1 it is a retreacable loop in the piecewise smoothly immersed class. In particular, $\gamma$ is not a whisker and by Tlas result we have the same conclusion as before.
@@
$\int_{t_{1}<\ldots<t_{q}}A(\gamma^{\prime}(t_{1}))\,\ldots\,A(\gamma^{\prime}(t_{1}))\,dt_{1}\ldots dt_{q}$
@@ Acknowledgments
We would like to thank Pablo Lessa for putting our attention on the Chen signature map.

### v4

@@ 5.1. Main lemma
If $\gamma\in\Omega^{\,psi}(\mathbb{R}^{n})$ is not trivial in ${\mathcal{E}}^{\,psi,ret}(\mathbb{R}^{n})$, then there are compactly supported one forms $\omega_{1},\ldots,\omega_{q}$ such that
@@
Since $SL(2,\,\mathbb{R})$ is simple, then Theorem 2.1 applies, that is, there is a connection $A$ in the bundle $\mathbb{R}^{n}\times SL(2,\mathbb{R})$, such that its holonomy $H_{A}(\gamma)$ is not trivial. This can also be deduced from Tlas result [Tl] and Proposition 1.1 as follows: since every whisker in the piecewise smooth class is a thin loop (see Remark $4$ in [Me]), by Proposition 1.1 it is a retraceable loop in the piecewise smoothly immersed class. In particular, $\gamma$ is not a whisker and by Tlas result we have the same conclusion as before.
@@
$\int_{t_{1}<\ldots<t_{q}}A(\gamma^{\prime}(t_{1}))\,\ldots\,A(\gamma^{\prime}(t_{q}))\,dt_{1}\ldots dt_{q}$
@@ Acknowledgments
We would like to thank Pablo Lessa for putting our attention on the Chen signature map and to Mauricio Velasco for his suggestions for improving the manuscript.